Match the descriptions to the steps used in solving the linear equation 3(3x−8)=4x+6.
1
Expert's answer
2020-09-28T15:20:59-0400

First of all expand the brackets on the left.

We get: 9x - 24 = 4x + 6

Move everything to the left part

Get: 9x - 4x - 24 - 6 = 0

9x - 4x = 5x

Get: 5x - 24 - 6 = 0

Move the numerical part to the right

We get: 5x = 24 + 6

24 + 6 = 30

Get: 5x = 30

Divide the right and left parts by 5

We get: x = 6

It is answer to this question